OpenStudy (anonymous):

Given that f(x) =-2xh(x) h(5) = -1 and h'(5)=-3 find f'(5)

OpenStudy (bahrom7893):

f(x) = -2x * h(x) f'(x) = -2x * h'(x) + (-2) * h(x)

OpenStudy (bahrom7893):

f'(5) = -2*5-h'(5) + (-2) * h(5)

OpenStudy (bahrom7893):

f'(5) = -10-(-3)+(-2) *(-1)

OpenStudy (bahrom7893):

f'(5) = -10+3+2=-5

OpenStudy (anonymous):

sorry mate that isn't one of the options

OpenStudy (anonymous):

not even close :-)

OpenStudy (bahrom7893):

I see my mistake..

OpenStudy (anonymous):

sweet, I am more concerned with how to understand this one

OpenStudy (bahrom7893):

-2*5*h'(5) + (-2) * h(5) f'(5) = -10*(-3)+(-2) *(-1) f'(5) = 30 + 2 = 32

OpenStudy (bahrom7893):

is that better? haha

OpenStudy (anonymous):

f(x) = -2xh(x) h(5) = -1 h'(5) = -3 f'(x) = -2x[h'(x)] + -2(h(x)) f'(5) = -2(5)(h'(5)) + -2(h(5)) = -10(-3) + -2(-1) = 30 + 2 = 32

OpenStudy (anonymous):

I agree.

OpenStudy (bahrom7893):

I don't know where that - came from haha..

OpenStudy (anonymous):

yes, much better. Im going through it to see where I went wrong _ thanks!!

OpenStudy (anonymous):

did you use the chain rule?

OpenStudy (anonymous):

pls explain if possible would really help

OpenStudy (anonymous):

why did you plug in the h's like you did?

OpenStudy (bahrom7893):

okay here we go: first use the chain rule: f'(x) = -2x * h'(x) + (-2) * h(x)

OpenStudy (anonymous):

Actually that's the product rule. I was wrong about the chain rule earlier.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

But what you have written is otherwise correct.

OpenStudy (bahrom7893):

*oh sorry it is.. i never remember the names..haha

OpenStudy (anonymous):

can I write the f(x) like this f(x) = (-2x)((h(x)) ?

OpenStudy (anonymous):

Certainly.

OpenStudy (bahrom7893):

yea..

OpenStudy (bahrom7893):

then use the product rule..

OpenStudy (anonymous):

perfect! now I see the product rule instead of me futzing around with the chain rule. Thanks a ton!!!

OpenStudy (bahrom7893):

sorry =P

OpenStudy (anonymous):

I was doing it loooong before I put the question up - def my fault